# Eliciting In-context Retrieval and Reasoning for Long-context Large Language Models

### v2

@@ Abstract
Recent advancements in long-context language models (LCLMs) promise to transform Retrieval-Augmented Generation (RAG) by simplifying pipelines. With their expanded context windows, LCLMs can process entire knowledge bases and perform retrieval and reasoning directly – a capability we define as In-Context Retrieval and Reasoning (ICR2). However, existing benchmarks like LOFT often overestimate LCLM performance by providing overly simplified contexts. To address this, we introduce ICR2, a benchmark that evaluates LCLMs in more realistic scenarios by including confounding passages retrieved with strong retrievers. We then propose three methods to enhance LCLM performance: (1) retrieve-then-generate fine-tuning, (2) retrieval-attention-probing, which uses attention heads to filter and de-noise long contexts during decoding, and (3) joint retrieval head training alongside the generation head. Our evaluation of five well-known LCLMs on LOFT and ICR2 demonstrates significant gains with our best approach applied to Mistral-7B: +17 and +15 points by Exact Match on LOFT, and +13 and +2 points on ICR2, compared to vanilla RAG and supervised fine-tuning, respectively. It even outperforms GPT-4-Turbo on most tasks despite being a much smaller model.
@@ 1 Introduction
The ability of large language models to process long contexts has significantly expanded their applicability across various domains, including book-level information retrieval (Ding et al., 2024; Jin et al., 2024), summarization (Kim et al., 2024; Saxena and Keller, 2024; Qiu et al., 2023), and question answering (Liu et al., 2024; Wang et al., 2024a). They also enable more complex tasks, such as agent trajectory modeling and planning (Zhao et al., 2024; Zhang et al., 2024b), video captioning (Xue et al., 2024; Zhang et al., 2023), and text-to-video generation (Wang et al., 2024b; Lin et al., 2023). Recent advancements in long-context language models (LCLMs) hold particular promise for reshaping the Retrieval-Augmented Generation (RAG) paradigm (Lee et al., 2024; Li et al., 2024). With their expanded context windows, LCLMs reduce reliance on complex pipelines required by previous context-length limitations and simplify knowledge updates by allowing context modifications. For instance, LCLMs can accommodate entire knowledge bases within their context windows, effectively serving as working memory to be used for new queries.
@@ 2 Related Work
Long-context large language models (LCLMs) have garnered significant attention for their ability to process extended sequences. Models like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020) introduced sparse attention mechanisms to efficiently handle long context. Scaling efforts, exemplified by GPT-4 (Achiam et al., 2023), underscore the importance of expanding context windows for the long-context tasks. Data engineering methods (Fu et al., 2024; Xiong et al., 2024; Jin et al., 2024), expanding positional encoding (Ding et al., 2024), and the parameter-efficient fine-tuning (Chen et al., 2024) have also been proven to be effective.
@@ 3.2 ICR2 Benchmark
Table 2 compares statistics between LOFT and our benchmark ICR2. An example of an ICR2 instance is provided in Table 1, and the rag_prompt template used is detailed in Table 8 in Appendix A.
Task #CTX #Tokens #Prov
LOFT NQ 215 28,911 1
HPQA 276 28,924 1.98
MUSIQ. 207 28,806 1.67
ICR2 NQ 202 20,441 1
HPQA 202 20,818 2.09
FEVER 202 21,136 1
WoW 202 21,068 1
@@ 4.1 Retrieve-then-generate Fine-tuning
Compared to the standard supervised fine-tuning where a model is asked to generate a Direct Answer (DA) to a given query, our first proposal, retrieve-then-generate, is a two-step process: the model first retrieves relevant passages from context and then generates the final answer based on the context and retrieved passages, all in one decoding pass. Formally, we train an LCLM to optimize the objective,
@@
We implement two variants of retrieve-then-generate fine-tuning, both using the special tokens <RETRIEVAL> and </RETRIEVAL> to delineate the retrieval step. In the first, Retrieve-Then-Answer (RTA), the model copies relevant passages verbatim during retrieval. In the second, Cite-Context-ID (CCI), the model generates only the IDs of the relevant passages. Figure 3 illustrates the templates for these variations.
@@ 4.3 Joint Retrieval Head Training
Figure 4: A retrieval head is asdded to predict the Top-K context for generating final responses. The head is jointly optimized with the generation head.
@@
We then set the binary mask $M$ accordingly to filter the original context $C$ down to $C^{*}$, which is then used to prompt the LCLM to produce the final response.
@@ 5.1 Setup
Baselines. We compare our approaches with the baselines in Sec. 3.3: Vanilla RAG, Closed-book, and Oracle RAG. We include the direct-answer SFT (SFT-DA) which follows Zhang et al. by concatenating the confounders with gold documents for SFT, and our methods in Sec. 4 include two retrieve-then-generate SFT variants — Retrieve-then-Answer (SFT-RTA), and Cite-Context-ID (SFT-CCI) — and the joint retrieval head training (RetHead) and retrieval attention probing (RAP).
@@ 5.2 Main Results
LOFT ICR2
NQ HPQA MUSI. NQ HPQA FEV WoW
@@
Table 4: Main results on LOFT (Lee et al., 2024) and ICR2 for our proposed methods applied on Mistral-2-7B. We also report the GPT-4 performance in the top panel. We highlight the improved and worsen performances compared with Vanilla RAG, and bold the best method.
@@
Among the SFT variants, SFT-RTA in general outperforms the others with an average improvement of 2%. Specifically on LOFT benchmark, both SFT-RTA and SFT-CCI outperform SFT-DA with 6% and 2% improvement on average, respectively, indicating the retrieve-then-generate strategy helps. On ICR2, however, all SFT variants perform the same. This demonstrates that ICR2 is a more discriminative benchmark than LOFT.
We apply RAP to all SFT models. On average, RAP enhances the models significantly, with SFT-DA + RAP improving by 6% and SFT-RTA + RAP by 8%. The best-performing approach, SFT-RTA + RAP, achieves notable gains on the challenging ICR2 benchmark, with improvements of 3%, 4%, 1%, and 1% on NaturalQuestions, HotpotQA, FEVER, and WoW, respectively. It also achieves top performance on 5 out of the 7 tasks, demonstrating its superiority.Remarkably, it achieves comparable performance with the state-of-the-art GPT-4-Turbo on LOFT and ICR2 while using a much smaller model. Finally, RAP decoding is more effective for SFT models than with the original model, as SFT better activates retrieval-specific attention heads for the approach (see Sec. 6.3).
@@ 5.3 In-Context Retrieval Performance
LCLM Methods LOFT ICR2 Avg.-All
NQ HPQA MUSI. Avg. NQ HPQA FEV. WoW Avg.
Mistral-2-7B SFT-RTA 0.81 0.83 0.37 0.67 0.58 0.78 0.64 0.43 0.61 0.63
SFT-CCI 0.65 0.61 0.26 0.51 0.63 0.73 0.69 0.49 0.64 0.58
@@
RetHead $w/\mathcal{L}_{gen}$ 0.00 0.00 0.00 0.02 0.01 -
RetHead $w/\mathcal{L}_{ret}$ 0.33 0.49 0.51 0.49 0.46 -
@@ 6.2 Blocking Context Attention in Retrieve-then-generate Model
To verify if models actually learn to generate the final responses only from the retrieval predictions in our retrieve-then-generate methods (Sec. 4.1), additional experiments were performed where we block attention to be paid onto the context beyond the retrieval predictions at the generation step, i.e.,
$p(y\mid q,c)=\sum_{z_{i}\in Z}p(y\mid q,z_{i})p(z_{i}\mid q,c),$ (6)
Table 6.2 reports the impact of the blocking on the SFT-RTA variant. We observe the performance essentially stays the same, indicating the model indeed learned to generate largely on the retrieved passages, ignoring the original context.
@@ 6.4 Retrieval Attention Probing
Figure 7 shows the high-level results on both LOFT and ICR2 (Appendix B has more details). As expected, we observe in the middle column that an increasing $M$ or $Q$ will increase the recall rate as the resulting larger pools of selected passages will more likely include the relevant ones. We also find that increasing $M$ as opposed to $Q$ is more effective in improving the recall rate as certain task such as HotpotQA requires multi-hop retrieval. However, the increased recall rate does not always translate to a better performance (left column) or RTA’s recall rate , as a higher $M$ or $Q$ may also introduce more confounders. This suggests the further reduction in confounding effects is still an opening future work.
@@ 7 Conclusion
In this paper, we introduce ICR2, a new benchmark designed as a more realistic and discriminative benchmark for evaluating LCLM’s abilities in in-context retrieval and reasoning. Our findings highlight the limitations for the current models. Additionally, we propose three methods —retrieve-then-generate fine-tuning, retrieval attention probing, and joint retrieval head training — to enhance models, achieving the results comparable to GPT-4, but with a smaller model footprint.
@@
Appendix A Prompt Template for the Retrieval-augmented Generation
@@
Appendix B Detailed Results for RAP Decoding
We show the detailed results for each task of LOFT and ICR2 in Figure 8 and 9.
@@ Appendix B Detailed Results for RAP Decoding
Appendix C Details of Experiment Setup
C.1 Training Details
@@ C.2.1 RAP Hyperparameter Settings
Appendix D Effect of Retrieval Delineation

### v3

@@ Abstract
Recent advancements in long-context language models (LCLMs) promise to transform Retrieval-Augmented Generation (RAG) by simplifying pipelines. With their expanded context windows, LCLMs can process entire knowledge bases and perform retrieval and reasoning directly – a capability we define as In-Context Retrieval and Reasoning (ICR2). However, existing benchmarks like LOFT often overestimate LCLM performance by providing overly simplified contexts. To address this, we introduce ICR2, a benchmark that evaluates LCLMs in more realistic scenarios by including confounding passages retrieved with strong retrievers. We then propose three methods to enhance LCLM performance: (1) retrieve-then-generate fine-tuning, (2) retrieval-attention-probing, which uses attention heads to filter and de-noise long contexts during decoding, and (3) joint retrieval head training alongside the generation head. Our evaluation of five well-known LCLMs on LOFT and ICR2 demonstrates significant gains with our best approach applied to Mistral-7B: +17 and +15 points by Exact Match on LOFT, and +13 and +2 points on ICR2, compared to vanilla RAG and supervised fine-tuning, respectively. It even outperforms GPT-4-Turbo on most tasks despite being a much smaller model.  [^1] Our code and datasets are available at https://github.com/apple/ml-icr2.
@@ 1 Introduction
The ability of large language models to process long contexts has significantly expanded their applicability across various domains, including book-level information retrieval (Ding et al., 2024; Jin et al., 2024), summarization (Kim et al., 2024; Saxena and Keller, 2024; Qiu et al., 2023), and question answering (Liu et al., 2024; Wang et al., 2024a). They also enable more complex tasks, such as agent trajectory modeling and planning (Zhao et al., 2024; Zhang et al., 2024b), video captioning (Xue et al., 2024; Zhang et al., 2023), and text-to-video generation (Wang et al., 2024b; Lin et al., 2023). Recent advancements in long-context language models (LCLMs) hold particular promise for reshaping the Retrieval-Augmented Generation (RAG) paradigm (Lee et al., 2024; Li et al., 2024). With their expanded context windows, LCLMs reduce reliance on complex pipelines required by context-length limitations and simplify knowledge updates by allowing context modifications. For instance, LCLMs can accommodate entire knowledge bases within the context windows, effectively serving as working memory for new queries.
@@ 2 Related Work
Long-context large language models (LCLMs) have garnered significant attention for their ability to process extended sequences. Models like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020) introduced sparse attention mechanisms to efficiently handle long context. Scaling efforts, exemplified by GPT-4 (Achiam et al., 2023), underscore the importance of expanding context windows for the long-context tasks. Data engineering methods (Fu et al., 2024; Xiong et al., 2024; Jin et al., 2024), expanding positional encoding (Ding et al., 2024), and the parameter-efficient fine-tuning (Chen et al., 2024) have been effective.
@@ 3.2 ICR2 Benchmark
Table 7 compares statistics between LOFT and our benchmark ICR2. An example of an ICR2 instance is provided in Table 1, and the rag_prompt template used is detailed in Table 8 in Appendix B.
@@ 4.1 Retrieve-then-generate Fine-tuning
Compared to the standard supervised fine-tuning where a model is asked to generate a Direct Answer (DA) to a given query, our first proposal, retrieve-then-generate, is a two-step process: the model first retrieves relevant passages from context and then generates the final answer based on the context and retrieved passages, all in one decoding pass. Formally, we train an LCLM to optimize,
@@
We use two variants of retrieve-then-generate fine-tuning, both using the special tokens <RETRIEVAL> and </RETRIEVAL> to delineate the retrieval step. In the first, Retrieve-Then-Answer (RTA), the model copies relevant passages. In the second, Cite-Context-ID (CCI), the model generates only the IDs of the relevant passages. Figure 8 shows the templates for all variations.
@@ 4.3 Joint Retrieval Head Training
We then set the binary mask $M$ for filtering the original context $C$ down to $C^{*}$, which is then used to prompt the LCLM to produce the final response.
@@ 5 Experiment
Baselines. We compare our approaches with the baselines in Sec. 3.3: Vanilla RAG, Closed-book, and Oracle RAG. We include the traditional ranking pipeline where we use TinyBERT (Jiao et al., 2020) and MiniLM (Wang et al., 2020) fine-tuned on MS Marco passage retrieval corpus (Bajaj et al., 2016) to select the top-relevant passages for LCLMs. We report their retrieval performance in Appendix D. We include the direct-answer SFT (SFT-DA) which follows Zhang et al. by concatenating the confounders with gold documents for SFT, and our methods in Sec. 4 include two retrieve-then-generate SFT variants — Retrieve-then-Answer (SFT-RTA), and Cite-Context-ID (SFT-CCI) — and the joint retrieval head training (RetHead) and retrieval attention probing (RAP).
@@ 5.1 Main Results
Models & Methods LOFT ICR2
NQ HotpotQA MUSIQUE NQ HotpotQA FEVER WoW
@@
Table 3: Main results on LOFT (Lee et al., 2024) and ICR2 for our methods applied on Mistral-2-7B-Instruct (Jiang et al., 2023). We also report the GPT-4 performance in the top panel. We highlight the improved and worsen performances compared with Vanilla RAG, and bold the best method based on Mistral-7B, except the Oracle.
@@
We first apply RTA-style generation as a prompt-only method (RTA Prompting). The resulting performance gains, though modest, demonstrate the potential of decoupling the generation process into separate retrieval and generation stages. Among the SFT variants, SFT-RTA in general outperforms the others with an average improvement of 2%. Specifically on LOFT benchmark, both SFT-RTA and SFT-CCI outperform SFT-DA with 6% and 2% improvement on average, respectively, indicating the retrieve-then-generate strategy helps. On ICR2, however, all SFT variants perform the same. This demonstrates that ICR2 is a more discriminative benchmark than LOFT.
We apply RAP to all SFT models. On average, RAP enhances the models significantly, with SFT-DA + RAP improving by 6% and SFT-RTA + RAP by 8%. The best-performing approach, SFT-RTA + RAP, achieves notable gains on the challenging ICR2 benchmark, with improvements of 3%, 4%, 1%, and 1% on NaturalQuestions, HotpotQA, FEVER, and WoW, respectively. It also achieves top performance on 5 out of the 7 tasks, demonstrating its superiority.Remarkably, it achieves comparable performance with the state-of-the-art GPT-4-Turbo on LOFT and ICR2 while using a much smaller model. Finally, RAP decoding is more effective for SFT models than with the original model, as SFT better activates retrieval-specific attention heads for the approach (see Sec. 6.3). To verify the generalization of methods across various LLMs, we conduct experiments with LLaMA-3-Instruct (Dubey et al., 2024) in Appendix C: we observe the similar improvements for retrieve-then-generate training, and RAP decoding.
Compared to the traditional pipeline, LCLMs can effectively unify in-context retrieval and generation, particularly on challenging benchmarks such as ICR2. This highlights the advantage of an end-to-end approach, wherein LCLMs contextually select retrieved items based on both the query and the evolving generation.
@@ 5.2 In-Context Retrieval Performance
Methods LOFT ICR2
NQ HPQA MUSI. NQ HPQA FEV. WoW
SFT-RTA 0.81 0.83 0.37 0.58 0.78 0.64 0.43
SFT-CCI 0.65 0.61 0.26 0.63 0.73 0.69 0.49
@@
$w/\mathcal{L}_{gen}+\mathcal{L}_{ret}$ 0.01 0.01 0.01 0.36 0.44 0.55 0.51
$w/\mathcal{L}_{gen}$ 0.00 0.00 0.00 0.02
$w/\mathcal{L}_{ret}$ 0.33 0.49 0.51 0.49
@@ 6.2 Blocking Context Attention in Retrieve-then-generate Model
To verify if models actually learn to generate the final responses only from the retrieval predictions in our retrieve-then-generate methods (Sec. 4.1), additional experiments were performed where we block attention to be paid onto the context beyond the retrieval predictions at the generation step. Table 6.2 reports the impact of the blocking on the SFT-RTA variant. We observe the performance essentially stays the same, indicating the model indeed learned to generate largely on the retrieved passages, ignoring the original context.
@@ 6.4 Retrieval Attention Probing
Figure 5 shows the high-level results on both LOFT and ICR2 (Appendix E has more details). As expected, we observe in the middle column that an increasing $M$ or $Q$ will increase the recall rate as the resulting larger pools of selected passages will more likely include the relevant ones. We also find that increasing $M$ as opposed to $Q$ is more effective in improving the recall rate as certain task such as HotpotQA requires multi-hop retrieval. However, the increased recall rate does not always translate to a better performance (left column) or RTA’s recall rate , as a higher $M$ or $Q$ may also introduce more confounders. This suggests the further reduction in confounding effects is still an opening future work.
@@ 7 Conclusion
In this paper, we introduce ICR2, a new benchmark designed as a more discriminative benchmark for evaluating LCLMs in in-context retrieval and reasoning. Our findings highlight the limitations for the current models. We propose three methods —retrieve-then-generate fine-tuning, retrieval attention probing, and joint retrieval head training — to enhance models, achieving the results comparable to GPT-4 with a smaller model footprint.
@@
Appendix A Statistics for LOFT and ICR2
Task #CTX #Tokens #Prov
LOFT NQ 215 28,911 1
HPQA 276 28,924 1.98
MUSIQ. 207 28,806 1.67
ICR2 NQ 202 20,441 1
HPQA 202 20,818 2.09
FEVER 202 21,136 1
WoW 202 21,068 1
We present the comparison in dataset statistics between LOFT and ICR2 in Table 7.
Appendix B Prompt Template for the Retrieval-augmented Generation
@@
Appendix C Generalization to Other LCLMs
LLaMA-3-Instruct LOFT ICR2
NQ HPQA NQ HPQA
Closebook 0.41 0.19 0.44 0.19
RAG 0.45 0.56 0.53 0.45
RAG w/ TinyBERT 0.89 0.89 0.55 0.48
Oracle 0.81 0.77 0.80 0.74
SFT-DA 0.80 0.60 0.51 0.52
SFT-RTA 0.84 0.63 0.51 0.57
SFT-DA w/ RAP 0.83 0.72 0.61 0.60
Table 9: LLaMA-3’s performances on NQ and HotpotQA for LOFT and ICR2with our methods on NQ-LOFT and HPQA-LOFT.
Table 9 presents the performance of the LLaMA-3-Instruct-8B model across two benchmarks—NQ-LOFT and HPQA-LOFT—evaluated under LOFT and ICR2. The results highlight several key trends. First, the baseline Closebook performs poorly, particularly on the HPQA dataset, underscoring the need for external information. Incorporating retrieval via RAG offers significant improvements, especially on HPQA, where performance increases from 0.19 to 0.56. The Oracle setting represents the upper bound of retrieval quality, and shows the strongest performance overall.
We observe that adding a reranker such as TinyBERT to RAG further improves results under the LOFT setup, reaching 0.89 on both NQ and HPQA—approaching Oracle-level performance. Among supervised fine-tuning methods, SFT-DA and SFT-RTA provide moderate improvements over vanilla RAG. Notably, integrating our proposed RAP mechanism (e.g., in SFT-DA w/ RAP) leads to consistent gains, especially under ICR2. For example, SFT-DA w/ RAP achieves the best NQ score (0.61), validating the effectiveness of RAP in enhancing decision quality during retrieval and answer generation.
Appendix D Retrieval Performance for Re-ranking Strategies
Re-ranker Model K LOFT ICR2
NQ HPQA MUSI. NQ HPQA FEV WoW
TinyBERT 8 0.99 0.98 0.80 0.61 0.65 0.68 0.35
MiniLM 8 1.00 0.98 0.82 0.66 0.65 0.70 0.27
TinyBERT 32 0.99 0.99 0.81 0.89 0.82 0.87 0.55
MiniLM 32 1.00 1.00 0.90 0.86 0.81 0.89 0.43
TinyBERT 50 0.99 0.99 0.82 0.91 0.84 0.92 0.60
MiniLM 50 1.00 1.00 0.92 0.91 0.88 0.91 0.50
SFT-RTA w/ RAP / 0.95 0.99 1.00 0.90 0.95 0.97 0.92
Table 10: Retrieval performance comparison of re-ranker models across LOFT and ICR2benchmarks.
Table 10 presents a comprehensive evaluation of different re-ranking strategies applied to retrieval-augmented generation (RAG) across two benchmark suites: LOFT (comprising NQ, HPQA, and MUSI.) and ICR2(including NQ, HPQA, FEV, and WoW). We compare baseline RAG without re-ranking, re-ranking using lightweight models (TinyBERT and MiniLM) at varying retrieval depths $k\in\{8,32,50\}$, and our best-performing variant (SFT-RTA w/ RAP). Results show consistent improvements with increased $k$, and MiniLM-based re-rankers often slightly outperform TinyBERT counterparts. Notably, the SFT-RTA w/ RAP model yields the highest performance across most datasets, particularly excelling in more challenging ICR2domains. These findings suggest the strong retrieval capabilities of our SFT-RTA approach compared to the traditional re-ranking strategies.
Appendix E Detailed Results for RAP Decoding
We show the detailed results for each task of LOFT and ICR2 in Figure 6 and 7.
@@ Appendix E Detailed Results for RAP Decoding
Appendix F Details of Experiment Setup
F.1 Training Details
@@ F.2.1 RAP Hyperparameter Settings
Appendix G Prompt Templates for Supervised Fine-tuning
We visualize three templates used for supervised fine-tuning LCLMs to perform retrieve-then-generate generation in Figure 8.
Appendix H Visualization for Joint Retrieval Head Training
We visualize the architecture for adding a retrieval head jointly trained with language generation in Figure 9 for easing the understanding of the approach.
Figure 9: A retrieval head is added to predict the Top-K context for generating final responses. The head is jointly optimized with the generation head.
Appendix I Effect of Retrieval Delineation